Compute to get the equilibrium price and quantity.


1) Qd= 30-5P


2) Qs= 0+5P


1
Expert's answer
2022-01-18T15:12:29-0500

Solution:

A.). At equilibrium: Qd = Qs

30 – 5P = 0 + 5P

30 – 0 = 5P + 5P

30 = 10P

P = 3

Equilibrium price = 3


Substitute to derive equilibrium quantity:

Qd = 30 – 5P = 30 – 5(3) = 30 – 15 = 15

Qd = 15

Equilibrium quantity = 15